Compute each sum below. If applicable, write your answer as a fraction. 3 + 3(4) +3(4) + ... +3() = 1 미 X 3 ?
Answer #1

1) a = 3 , r = 4 , 8

sum(S) = a(r^n-1)/(r-1)

S = 3(4^8-1)/3 = 64435 [Answer]

2)

a = -1/2 , r = -1/2 , n =9

sum = (-1/2)((-1/2)^9-1)/(-3/2) = -171/512 [Answer]